OpenStudy (anonymous):

how do you simplify 3/7 - 4/m

OpenStudy (lgbasallote):

\[\huge \frac 3{7 - \frac 4m}\] ???

OpenStudy (anonymous):



OpenStudy (anonymous):

\[A. \frac{ 3m-28 }{ 7m } B. -\frac{ 1 }{ 7m} C. -\frac{ 25 }{7m} D. -\frac{ 25 }{ 7}\]

OpenStudy (anonymous):

how do i simiplify it though?

OpenStudy (anonymous):

multiply not multiplay lol

OpenStudy (anonymous):

You must get a common denominator.. My sugestion multiplay the left side by m/m and the right side by 7/7 then you can just subtract the numerators

OpenStudy (anonymous):

Answer A is the answer to the question as reiterated by Abirla above. We get it in the following manner\[\frac{3}{7}-\frac{4}{m}=\frac{3}{7}\frac{m}{m}-\frac{4}{m}\frac{7}{7}=\frac{3m}{7m}-\frac{28}{7m}=\frac{3m-28}{7m}\]

OpenStudy (anonymous):

The answer is (a). You need to find a common denominator. Multiply (3/7) by m/m which is essentially 1. Also multiply 4/m by 7/7. Now you can do the subtraction to get (a)

OpenStudy (anonymous):

As Chyme said just above, multiply each term by one in a form that will give a common denominator (in this case, 7m).

OpenStudy (anonymous):

exact what i was thinking' the denominator is 7m

hero (hero):

Ha, I like multiplay better than multiply